# MotionFix: Text-Driven 3D Human Motion Editing

### v1

@@ 4 Text-Driven Motion Editing Diffusion Model
Figure 3: Models overview: (left) We illustrate our TMED model during training. We noise the target motion for $t$ steps, and the transformer model is trained to denoise it back by one step. The conditions – text and source motion – are appended to the input. CLIP backbone is frozen, while components denoted in pink are learned during training. At test time, the iterative diffusion process is initialized from random noise instead of the noised target. (right) Our MDM-BP baseline is repurposed from a pretrained text-to-motion generation model to be used only at test time for motion editing. The model is initialized from random noise and the body parts not to be edited according to GPT are copied from the source motion.
@@
We introduce TMED, a text-driven motion editing diffusion model. Given a short 3D human motion, a textual instruction describing a modification, and a noise vector to enable randomness, the model generates an edited motion. Similar tasks have been addressed in the image domain for text-based image editing [5], from which we take inspiration for our model design. We further build on motion diffusion model (MDM) [48] that takes only text as input and generates a motion. In contrast, our model has an additional condition on the source, thus requires a different training dataset (as described in Section 3). In the following, we present the components of our TMED model.
4.1 3D Human Motion Representation
@@
We use a sequence of SMPL [31] body parameters to represent a human motion. SMPL is a linear function that maps the shape, and pose parameters of $J$ joints, along with the global body translation and orientation, to a 3D mesh. The joint positions, $J_{p}$, can be obtained from vertices via the learned SMPL joint regressor. Following previous work [36] that discards the shape parameters, we set the shape parameters to zero (mean shape), since motion is parameterized primarily by pose parameters.
@@
Various alternative representations have been used based on joint positions with respect to the local coordinate system of the body [17, 21, 46]. Unlike prior works [48, 17, 51] that fit SMPL bodies to skeleton generations, we aim to enable direct regression of SMPL parameters, bypassing the need of a costly post-processing optimization [4], and thus making our method ready to use for animation frameworks.
@@
A common approach for representing SMPL pose parameters within a learning framework is to employ 6D rotations [53], and to apply first-frame canonicalization for motions [36, 2]. Similarly, we canonicalize our motions prior to training, so that all face the same direction in the first frame and have the same initial global position. Inspired by [21, 38], we represent the global body translation as differences between consecutive frames. Supervising with such relative translations helps the denoiser to generate better trajectories, as we observed unsmooth generations when using the absolute translation. Similar to STMC [38], we factor out the $z$-rotation from the pelvis orientation and separately represent the global orientation as the $xy$-orientation and the $z$-orientation as the differences between rotations in consecutive frames (resulting in $12$ features, i.e., 6D representation for $xy$ and $z$). We represent the body pose with 6D rotations [53]. Similar to [36], we exclude the hand joints as they mostly do not move in the datasets we use. We additionally append the local joint positions after removing $z$-rotation of the body [21, 38] (resulting in 192 dimensional features with $6\times 21$ for rotations and $22\times 3$ for joints including the root joint). Thus, each motion frame has a dimension $d_{p}=207$, consisting of $3$ features for the global translation, $12$ for the global orientation, and $192$ for the body pose. The motion is represented as a sequence of the pose representations. During training, all features are normalized according to their mean and variance over the training set.
4.2 Conditional Diffusion Model
@@
To learn TMED, we use our new training data, where each data sample comprises a source motion $M_{S}$, target motion $M_{T}$, and a language instruction $L$. We train a conditional diffusion model that learns to edit the source motion with respect to the instruction. We design a model similar to that of InstructPix2Pix [5], where the generation from a random noise vector is conditioned on two further inputs $L$ and $M_{S}$. Here, instead of a sequence of image patch tokens, the motion modality is represented as a variable-length sequence of motion frames. The noised target motion, the text condition $L$, and the source motion condition $M_{S}$ are all fed as input to the denoiser at every diffusion step.
@@
Diffusion models [45] learn to gradually turn random noise into a sample from a data distribution by a sequence of denoising autoencoders. This is achieved by a diffusion process that adds noise $\epsilon_{t}$ to an input signal, $M_{T}$. We denote the noise level added to the input signal by using $t$, the diffusion timestep, as a superscript. This produces a diffused sample, ${M}^{t}_{T}$. The amount of noise added at timestep $t=1,\dots,N$ is defined a-priori through a noise schedule. We train a denoiser network, to reverse this process given the timestep $t$, the instruction $L$, the noised target motion ${M}^{t}_{T}$ and the source motion $M_{S}$. As supervision, the output of the denoiser network $\tilde{M}^{t}_{T}$ is compared against the ground-truth denoised target motion $M_{T}$. Our model is therefore trained to minimize:
@@
The architecture overview is illustrated in Figure 3 (left). Our model consists of multiple encoders for each input modality ($E_{T}$ for timestep, $E_{L}$ for text, and $E_{M}$ for motion) and a transformer encoder $D$ that operates on all inputs. The timestep $t$ is encoded via $E_{T}$ similar to MDM [48], by first converting into a sinusoidal positional embedding, and then projecting through a feed-forward network (consisting of two linear layers with a SiLU activation [13] in between). As in [48], we use the CLIP [42] text encoder for $E_{L}$. We pass the source and noised target motions through a linear layer ($E_{M}$), shared across frames, and obtain $M^{enc}_{S}=E_{M}(M_{S})$ and ${M^{t}_{T}}^{enc}=E_{M}({M}^{t}_{T})$. Given the variable duration of source and target motions, we add a learnable separation token SEP in between [11] when appending them (so that the information on when the target motion ends and the source motion starts is communicated to the transformer). Once all encoded inputs have the same feature dimensionality $d$, they are combined into a single sequence to be fed to the transformer, as shown in Figure 3, and sinusoidal positional embeddings are subsequently added. During training, to enable classifier-free guidance, the source motion condition is randomly dropped 5% of the time, the text condition 5%, both conditions together 5%, and all the inputs are used 85% of the time. For sampling from a diffusion model with two conditions, we apply classifier-free guidance with respect to two conditions: the input motion $M_{S}$ and the text instruction $L$. We introduce separate guidance scales $s_{M_{S}}$ and $s_{L}$ that allow adjusting the influence of each conditioning.
In the following, for simplicity, we abuse the notation by dropping the timestep subscripts when deriving the sampling process. Our generative model, TMED, learns the probability distribution over the target motions, $M_{T}$ conditioned on the source motions and text condition, $P(M_{T}\mid M_{S},L)$. Expanding this conditional probability gives:
$\begin{split}P(M_{T}\mid M_{S},L)&=\frac{P(M_{T},M_{S},L)}{P(L,M_{S})}\\ &=\frac{P(L\mid M_{S},M_{T})P(M_{S}\mid M_{T})P(M_{T})}{P(L,M_{S})}.\end{split}$ (2)
As in the original diffusion, we formulate this as a score function optimization problem by first taking the logarithm of Eq.(2),
$\begin{split}\log(P(M_{T}\mid M_{S},L))&=\log(P(L\mid M_{S},M_{T}))\\ &\quad+\log(P(M_{S}\mid M_{T}))\\ &\quad+\log(P(M_{T}))\\ &\quad-\log(P(L,M_{S})),\end{split}$ (3)
and then the derivative with respect to the input of Eq.(3) gives the score estimate $\tilde{e}_{\theta}(M_{T},s_{M_{S}},s_{L})$, which is learned under classifier-free guidance:
@@
We further ablate the guidance scales which control the generation at test time in Section 5, and show that as we increase the motion guidance, the generations tend to move towards the source motion as expected.
Implementation details. All models are trained for $1000$ epochs using cosine noise schedule with DDPM scheduler [20]. We use $N=300$ diffusion timesteps, as we find this is a good compromise between speed and quality. The guidance scales are chosen for each model based on their best performance in the validation set of MotionFix. We follow the same process for training MDM [48] baselines described in the next section. In terms of architectural details, the dimensionality of the embeddings before inputting to the transformer is $d=512$. We use a pre-trained and frozen CLIP [42] with all 77 token outputs of the ViT-B/32 backbone [12] as our text encoder $E_{L}$. We use the text masks from $E_{L}$ to mask the padded area of the text inputs. The motion encoder $E_{M}$ that precedes the transformer is a simple linear projection with dimensionality $d_{p}\times d$, where the feature dimension of each motion frame is $d_{p}=207$ (as described in Section 4.1).
5 Experiments
@@
R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ AvgR $\downarrow$ R@1 $\uparrow$ R@2 $\uparrow$ R@3 $\uparrow$ AvgR $\downarrow$
GT n/a n/a 100.0 100.0 100.0 1.02 73.15 84.09 89.49 2.09
MDM HumanML3D 21.45 31.96 38.35 9.64 17.61 26.56 34.52 10.33
MDMS HumanML3D ✓, init 21.31 32.39 40.20 9.39 17.76 27.70 35.80 9.66
MDM-BPS HumanML3D ✓, init&BP 54.97 69.46 75.00 3.77 74.43 81.96 85.37 2.74
MDM-BP HumanML3D ✓, BP 55.26 69.32 76.70 3.93 74.72 81.68 85.23 2.91
TMED MotionFix ✓, condition 62.22 74.15 81.96 3.01 80.82 89.77 93.75 1.62
Table 2: Results on the MotionFix benchmark: We first evaluate several variants of our text-to-motion synthesis baseline (MDM) on the motion editing task. Subscript $S$ denotes models that denoise the source motion initialization (init) instead of starting the diffusion from noise. BP indicates GPT-based body part labeling described in Section 4 to mask the source body parts which are kept unchanged during diffusion. Our model TMED effectively learns how to utilize the source motion conditioning, thanks to the MotionFix training data. See text for detailed comments.
@@
Table 3: The effect of training data size in MotionFix: We observe significant performance improvement as we increase the amount of training data.
5.1 Evaluation Metrics
@@
Similar to text-to-motion synthesis, distance-based metrics to evaluate motion generation quality is problematic due to multiple plausible ground-truth motions for a given text. Prior work has extensively used text-to-motion retrieval metrics for evaluating text-to-motion synthesis [18, 48], by training a text-motion contrastive model and using its features. To evaluate motion editing, we introduce motion-to-motion retrieval metrics. Given a generated motion, we measure how well the source (generated-to-source retrieval) or the target motion (generated-to-target retrieval) can be retrieved. We use TMR [37] as the feature extractor, but train it ourselves to support our feature representation, using the same regime as in the original paper with HumanML3D data [17]. We report standard metrics, R@1, R@2, R@3 and AvgR using a gallery size of $32$ randomly sampled batches for retrieval. Recall at rank $k$ (R@k) computes the percentage of times the correct motion is among the top $k$ results. Note that we fix the batches so there is no randomness across evaluations. The performance is averaged across batches. We report results on the full test set as gallery in Tables A.2 and A.3 of the appendix, where the same conclusions hold. While the main performance measure is according to generated-to-target retrieval, we also monitor how close our generations remain to the source. As indicative values, we provide ground-truth (GT) values for the latter. We provide additional measures (FID, L2) and perceptual studies in Sections B and C of the appendix, respectively, which further confirm the results of our proposed retrieval-based benchmarking.
5.2 Comparison to Baselines
@@
We report our main results in Table 2. We compare the performance of TMED trained on our MotionFix triplets against several baselines trained on the larger, HumanML3D text-motion pairs [17]. We build our baselines by training MDM [48] with our human motion representation and by repurposing this model for motion editing, described next.
@@
We first introduce two simple baselines: (a) MDM that purely uses the edit text as input to the text-to-motion generation (i.e., without a source motion), and (b) MDMS that additionally uses the source motion as input instead of noise during inference. For the latter, we also investigated reducing the number of diffusion steps when initializing from source; however, we observed performance drops and therefore kept the full $300$ diffusion steps. Inspired by [3], we design two additional strong baselines (MDM-BPS and MDM-BP), that are based on body part labels extracted by querying GPT with the edit texts. We automatically detect body parts which are irrelevant to the text and keep them constant via masking. We again initialize the diffusion process either from the source motion (MDM-BPS) or from noise (MDM-BP) for the body parts that need to change according to the GPT response. For more details on the query and example GPT outputs, we refer to Section D of the appendix.
@@
We first observe from Table 2 that, for all the baselines, initializing from noise performs better than initializing from source motion. Our strong baselines based on body-part detection (MDM-BP, MDM-BPS) clearly outperform the naive baselines. However, all baselines fall behind our TMED that successfully leverages the access to training triplets, and significantly outperforms alternatives. Moreover, MDM-BP, MDM-BPS are both strong baselines, but relying on GPT body part labels might not capture all edit types, such as the ones that require modifying the overall body. We demonstrate this further in our qualitative comparisons (Section 5.4).
5.3 Ablations

### v3

@@ 4. Text-Driven Motion Editing Diffusion Model
Figure 3. Models overview: (left) We illustrate our TMED model during training. We noise the target motion for $t$ steps, and the transformer model is trained to denoise it back by one step. The conditions – text and source motion – are appended to the input. CLIP backbone is frozen, while components denoted in pink are learned during training. At test time, the iterative diffusion process is initialized from random noise instead of the noised target. (right) Our MDM-BP baseline is repurposed from a pretrained text-to-motion generation model to be used only at test time for motion editing. The model is initialized from random noise and the body parts not to be edited according to GPT are copied from the source motion.
[…]
Model overview
@@
We introduce TMED, a text-driven motion editing diffusion model. Given a short 3D human motion, a textual instruction describing a modification, and a noise vector to enable randomness, the model generates an edited motion. Similar tasks have been addressed in the image domain for text-based image editing (Brooks et al., 2023), from which we take inspiration for our model design. We further build on motion diffusion model (MDM) (Tevet et al., 2023) that takes only text as input and generates a motion. In contrast, our model has an additional condition on the source, thus requires a different training dataset (as described in Section 3). In the following, we present the components of our TMED model.
4.1. 3D Human Motion Representation
@@
We use a sequence of SMPL (Loper et al., 2015) body parameters to represent a human motion. SMPL is a linear function that maps the shape, and pose parameters of $J$ joints, along with the global body translation and orientation, to a 3D mesh. The joint positions, $J_{p}$, can be obtained from vertices via the learned SMPL joint regressor. Following previous work (Petrovich et al., 2022) that discards the shape parameters, we set the shape parameters to zero (mean shape), since motion is parameterized primarily by pose parameters.
@@
Various alternative representations have been used based on joint positions with respect to the local coordinate system of the body (Guo et al., 2022a; Holden et al., 2016; Starke et al., 2019). Unlike prior works (Tevet et al., 2023; Guo et al., 2022a; Zhang et al., 2023a) that fit SMPL bodies to skeleton generations, we aim to enable direct regression of SMPL parameters, bypassing the need of a costly post-processing optimization (Bogo et al., 2016), and thus making our method ready to use for animation frameworks.
@@
A common approach for representing SMPL pose parameters within a learning framework is to employ 6D rotations (Zhou et al., 2019), and to apply first-frame canonicalization for motions (Petrovich et al., 2022; Athanasiou et al., 2022). Similarly, we canonicalize our motions prior to training, so that all face the same direction in the first frame and have the same initial global position. Inspired by (Holden et al., 2016; Petrovich et al., 2024), we represent the global body translation as differences between consecutive frames. Supervising with such relative translations helps the denoiser to generate better trajectories, as we observed unsmooth generations when using the absolute translation. Similar to STMC (Petrovich et al., 2024), we factor out the $z$-rotation from the pelvis orientation and separately represent the global orientation as the $xy$-orientation and the $z$-orientation as the differences between rotations in consecutive frames (resulting in $12$ features, i.e., 6D representation for $xy$ and $z$). We represent the body pose with 6D rotations (Zhou et al., 2019). Similar to (Petrovich et al., 2022), we exclude the hand joints as they mostly do not move in the datasets we use. We additionally append the local joint positions after removing $z$-rotation of the body (Holden et al., 2016; Petrovich et al., 2024) (resulting in 192 dimensional features with $6\times 21$ for rotations and $22\times 3$ for joints including the root joint). Thus, each motion frame has a dimension $d_{p}=207$, consisting of $3$ features for the global translation, $12$ for the global orientation, and $192$ for the body pose. The motion is represented as a sequence of the pose representations. During training, all features are normalized according to their mean and variance over the training set.
4.2. Conditional Diffusion Model
@@
To learn TMED, we use our new training data, where each data sample comprises a source motion $M_{S}$, target motion $M_{T}$, and a language instruction $L$. We train a conditional diffusion model that learns to edit the source motion with respect to the instruction. We design a model similar to that of InstructPix2Pix (Brooks et al., 2023), where the generation from a random noise vector is conditioned on two further inputs $L$ and $M_{S}$. Here, instead of a sequence of image patch tokens, the motion modality is represented as a variable-length sequence of motion frames. The noised target motion, the text condition $L$, and the source motion condition $M_{S}$ are all fed as input to the denoiser at every diffusion step.
@@
Diffusion models (Sohl-Dickstein et al., 2015) learn to gradually turn random noise into a sample from a data distribution by a sequence of denoising autoencoders. This is achieved by a diffusion process that adds noise $\epsilon_{t}$ to an input signal, $M_{T}$. We denote the noise level added to the input signal by using $t$, the diffusion timestep, as a superscript. This produces a diffused sample, ${M}^{t}_{T}$. The amount of noise added at timestep $t=1,\dots,N$ is defined a-priori through a noise schedule. We train a denoiser network, to reverse this process given the timestep $t$, the instruction $L$, the noised target motion ${M}^{t}_{T}$ and the source motion $M_{S}$. As supervision, the output of the denoiser network $\tilde{M}^{t}_{T}$ is compared against the ground-truth denoised target motion $M_{T}$. Our model is therefore trained to minimize:
@@
The architecture overview is illustrated in Figure 3 (left). Our model consists of multiple encoders for each input modality ($E_{T}$ for timestep, $E_{L}$ for text, and $E_{M}$ for motion) and a transformer encoder $D$ that operates on all inputs. The timestep $t$ is encoded via $E_{T}$ similar to MDM (Tevet et al., 2023), by first converting into a sinusoidal positional embedding, and then projecting through a feed-forward network (consisting of two linear layers with a SiLU activation (Elfwing et al., 2018) in between). As in (Tevet et al., 2023), we use the CLIP (Radford et al., 2021) text encoder for $E_{L}$. We pass the source and noised target motions through a linear layer ($E_{M}$), shared across frames, and obtain $M^{enc}_{S}=E_{M}(M_{S})$ and ${M^{t}_{T}}^{enc}=E_{M}({M}^{t}_{T})$. Given the variable duration of source and target motions, we add a learnable separation token SEP in between (Devlin et al., 2019) when appending them (so that the information on when the target motion ends and the source motion starts is communicated to the transformer). Once all encoded inputs have the same feature dimensionality $d$, they are combined into a single sequence to be fed to the transformer, as shown in Figure 3, and sinusoidal positional embeddings are subsequently added. During training, to enable classifier-free guidance, the source motion condition is randomly dropped 5% of the time, the text condition 5%, both conditions together 5%, and all the inputs are used 85% of the time. For sampling from a diffusion model with two conditions, we apply classifier-free guidance with respect to two conditions: the input motion $M_{S}$ and the text instruction $L$. We introduce separate guidance scales $s_{M_{S}}$ and $s_{L}$ that allow adjusting the influence of each conditioning.
For simplicity, we now abuse the notation by dropping the timestep subscripts when deriving the sampling process. Our generative model, TMED, learns the probability distribution over the target motions, $M_{T}$, conditioned on the source motions and text condition, $P(M_{T}\mid M_{S},L)$. Expanding this conditional probability gives:
(2) $\begin{split}P(M_{T}\mid M_{S},L)&=\frac{P(M_{T},M_{S},L)}{P(L,M_{S})}\\ &=\frac{P(L\mid M_{S},M_{T})P(M_{S}\mid M_{T})P(M_{T})}{P(L,M_{S})}.\end{split}$
As in the original diffusion, we formulate this as a score function optimization problem by first taking the logarithm of Eq.(2):
(3) $\begin{split}\log(P(M_{T}\mid M_{S},L))&=\log(P(L\mid M_{S},M_{T}))\\ &\quad+\log(P(M_{S}\mid M_{T}))\\ &\quad+\log(P(M_{T}))\\ &\quad-\log(P(L,M_{S})).\end{split}$
Then, the derivative with respect to the input of Eq.(3) gives the score estimate $\tilde{e}_{\theta}(M_{T},s_{M_{S}},s_{L})$, learned under classifier-free guidance:
@@
We further ablate the guidance scales which control the generation at test time in Section 5.
Implementation details. All models are trained for $1000$ epochs using cosine noise schedule with DDPM scheduler (Ho et al., 2020). We use $N=300$ diffusion timesteps, as we find this is a good compromise between speed and quality. The guidance scales are chosen for each model based on their best performance in the validation set of MotionFix ($S_{L}=2,s_{M_{S}}=2$). We follow the same process for training MDM (Tevet et al., 2023) baselines described in the next section. In terms of architectural details, the dimensionality of the embeddings before inputting to the transformer is $d=512$. We use a pre-trained and frozen CLIP (Radford et al., 2021) with all 77 token outputs of the ViT-B/32 backbone (Dosovitskiy et al., 2021) as our text encoder $E_{L}$. We use the text masks from $E_{L}$ to mask the padded area of the text inputs. The motion encoder $E_{M}$ that precedes the transformer is a simple linear projection with dimensionality $d_{p}\times d$, where the feature dimension of each motion frame is $d_{p}=207$ (as described in Section 4.1).
5. Experiments
@@
R@1 R@2 R@3 AvgR R@1 R@2 R@3 AvgR
GT n/a n/a 100.0 100.0 100.0 1.00 74.01 84.52 89.91 2.03
MDM HumanML3D 4.03 7.56 10.48 15.55 2.62 6.15 9.38 15.88
MDMS HumanML3D ✓, init 3.63 7.06 10.08 15.64 2.62 6.25 9.78 15.84
MDM-BPS HumanML3D ✓, init&BP 38.10 48.99 54.84 6.47 60.28 69.46 73.89 4.23
MDM-BP HumanML3D ✓, BP 39.10 50.09 54.84 6.46 61.28 69.55 73.99 4.21
TMED MotionFix ✓, condition 62.90 76.51 83.06 2.71 71.77 84.07 89.52 1.96
Table 2. Results on the MotionFix benchmark (test set): We first evaluate several variants of our text-to-motion synthesis baseline (MDM) on the motion editing task. Subscript $S$ denotes models that denoise the source motion initialization (init) instead of starting the diffusion from noise. BP indicates GPT-based body part labeling described in Section 4 to mask the source body parts which are kept unchanged during diffusion. Our model TMED effectively learns how to utilize the source motion conditioning, thanks to the MotionFix training data. See text for detailed comments.
@@
Table 3. Effect of training data size in MotionFix: We observe significant performance improvement as we increase the amount of training data.
5.1. Evaluation Metrics
@@
Similar to text-to-motion synthesis, distance-based metrics to evaluate motion generation quality is problematic due to multiple plausible ground-truth motions for a given text. Prior work has extensively used text-to-motion retrieval metrics for evaluating text-to-motion synthesis (Guo et al., 2022b; Tevet et al., 2023), by training a text-motion contrastive model and using its features. To evaluate motion editing, we introduce motion-to-motion retrieval metrics. Given a generated motion, we measure how well the source (generated-to-source retrieval) or the target motion (generated-to-target retrieval) can be retrieved. We use TMR (Petrovich et al., 2023) as the feature extractor, but train it ourselves to support our feature representation, using the same regime as in the original paper with HumanML3D data (Guo et al., 2022a). We report standard metrics, R@1, R@2, R@3 and AvgR using a gallery size of $32$ randomly sampled batches for retrieval from the test set. Recall at rank $k$ (R@k) computes the percentage of times the correct motion is among the top $k$ results. Note that we fix the batches so there is no randomness across evaluations. The performance is averaged across batches. We report results on the full test set as gallery in Tables A.2 and A.3 of the appendix, where the same conclusions hold. While the main performance measure is according to generated-to-target retrieval, we also monitor how close our generations remain to the source. As indicative values, we provide ground-truth (GT) values for the latter. We provide additional measures (FID, L2) and perceptual studies in Sections B and C of the appendix, respectively, which further confirm the results of our proposed retrieval-based benchmarking.
5.2. Comparison to Baselines
@@
We report our main results in Table 2. We compare the performance of TMED trained on our MotionFix triplets against several baselines trained on the larger, HumanML3D text-motion pairs (Guo et al., 2022a). We build our baselines by training MDM (Tevet et al., 2023) with our human motion representation and by repurposing this model for motion editing, described next.
@@
We first introduce two simple baselines: (a) MDM that purely uses the edit text as input to the text-to-motion generation (i.e., without a source motion), and (b) MDMS that additionally uses the source motion as input instead of noise during inference. For the latter, we also investigated reducing the number of diffusion steps when initializing from source; however, we observed performance drops and therefore kept the full $300$ diffusion steps. Inspired by (Athanasiou et al., 2023), we design two additional strong baselines (MDM-BPS and MDM-BP), that are based on body part labels extracted by querying GPT with the edit texts. We automatically detect body parts which are irrelevant to the text and keep them constant via masking. We again initialize the diffusion process either from the source motion (MDM-BPS) or from noise (MDM-BP) for the body parts that need to change according to the GPT response. For more details on the query and example GPT outputs, we refer to Section D of the appendix.
@@
We first observe from Table 2 that, for all the baselines, initializing from noise performs better than initializing from source motion. Our strong baselines based on body-part detection (MDM-BP, MDM-BPS) clearly outperform the naive baselines. However, all baselines fall behind our TMED that successfully leverages the access to training triplets, and significantly outperforms alternatives.
Moreover, MDM-BP, MDM-BPS are both strong baselines, but relying on GPT body part labels might not capture all edit types, such as the ones that require modifying the overall body. We demonstrate this further in our supplementary video from the project webpage and our qualitative comparisons (Section 5.4).
5.3. Ablations